acar that had stopped at the traffic light started at acceleration


of 2 m/s2.at the same time atruk overtook the car at aconstant


velocity 10 m/s                                                       


a) after how many meters will the car be beside the truk?


b)what will the velocity of the car be when it overtook by the


truk?

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Student

For car

initial velocity at t=0 is uc=0 and acceleration ac=2

For truck

initial velocity at t=0 is ut=10 and acceleration at=0

they will meet when Sc=St

uct + act2/2=utt + att2/2

0 + 2t2/2=10t + 0

t=10 sec

(a) Sc=act2/2=2*102/2 = 100m

(b) Vc=uc + act = 0 + 2*10=20 m/s

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..